# Multi-SpaCE: Multi-Objective Subsequence-based Sparse Counterfactual Explanations for Multivariate Time Series Classification

### v1

@@ 4.4 Evaluation metrics
  • Proximity: quantified with the $\ell_{2}$ distance between the original sample $x$ and the counterfactual $x^{\prime}$.
  • Sparsity: evaluated as the number of changes of the counterfactual, normalized by the total length of the time series: $\frac{||\mathbf{M}||_{0}}{L\times C}$. A lower score is preferred.
@@
We also compare the execution times and present them in Appendix D.
@@ 4.5 Results
In this section, we present the results in terms of validity, proximity, sparsity, plausibility, and contiguity. For all metrics (except for validity), all metrics reported are calculated exclusively on valid CFEs. The results are organized into tables specific to each metric, with separate subtables for univariate and multivariate datasets. The best performances for each dataset are highlighted in bold, while the second-best results are underlined. Additionally, we rank the methods for each dataset and metric and report their average rank to provide an intuitive performance comparison across different settings.
@@
AWR 0.99 0.92 0.24 1.0
BasicMotions 1.0 0.68 0.6 1.0
Cricket 0.99 0.9 - 1.0
Epilepsy 1.0 0.69 0.69 1.0
NATOPS 0.92 0.86 0.07 1.0
PEMS-SF 0.77 0.78 0.0 1.0
PenDigits 0.99 0.9 0.84 1.0
RacketSports 1.0 0.79 0.37 1.0
SR-SCP1 0.98 0.82 0.17 1.0
UWave 1.0 0.93 0.37 1.0
Average Rank 1.7 2.9 3.89 1.0
@@
Coffee 1.0 0.36 0.57 1.0 0.93 1.0
ECG200 1.0 0.04 0.07 0.5 0.53 1.0
FordA 1.0 0.48 0.48 0.97 0.57 1.0
Gunpoint 1.0 0.56 0.2 1.0 0.78 1.0
HandOutlines 1.0 0.84 0.52 0.52 0.68 1.0
ItalyPower 1.0 0.07 0.08 0.57 0.25 1.0
PPOC 1.0 0.7 0.08 0.16 0.43 1.0
Strawberry 1.0 0.98 0.7 0.82 0.87 1.0
CBF 1.0 - - 1.0 0.76 1.0
CinCECGTorso 1.0 - - 1.0 0.52 1.0
TwoPatterns 1.0 - - 1.0 0.99 1.0
ECG5000 1.0 - - 0.08 0.34 1.0
Plane 1.0 - - 0.99 0.87 1.0
FacesUCR 1.0 - - 0.95 0.95 1.0
NI-ECG2 1.0 - - 0.99 1.0 1.0
Average Rank 1.0 4.62 5.38 2.93 3.6 1.0
@@
Proximity results are presented in Table 3. While Multi-SpaCE does not directly optimize for proximity, it achieves competitive results. For multivariate datasets, Multi-SpaCE and DiscoX alternate between first and second place across datasets, with Multi-SpaCE achieving the best average rank by a small margin. In univariate datasets, Glacier is the top performer overall, frequently achieving the lowest proximity scores, followed by Multi-SpaCE ranking as the second-best method on average, and performing the best on most datasets where Glacier is not applicable. However, note that the number of valid counterfactuals generated by Glacier is often very low, thus reducing the relevance of its good proximity score.
@@
AWR 26.3 29.33 11.0 15.14
BasicMotions 103.36 89.51 68.75 62.91
Cricket 79.42 74.65 - 47.36
Epilepsy 20.11 14.43 11.63 9.44
NATOPS 7.87 14.82 6.2 7.95
PEMS-SF 4.48 10.55 - 5.04
PenDigits 86.32 90.82 81.31 45.11
RacketSports 67.76 59.71 32.31 41.65
SR-SCP1 764.11 815.47 673.94 428.47
UWave 30.68 23.39 15.63 14.42
Average Rank 3.1 3.3 1.62 1.5
@@
Coffee 1.21 0.22 12.3 1.59 3.44 1.0
ECG200 4.07 0.15 1.83 3.3 4.85 2.47
FordA 13.68 0.52 5.8 15.26 11.22 5.85
Gunpoint 2.47 0.48 3.93 2.99 4.09 2.44
HandOutlines 3.81 0.56 0.92 4.31 7.79 2.04
ItalyPower 1.6 0.22 1.43 1.53 2.03 1.33
PPOC 0.83 0.12 0.21 0.29 1.81 0.4
Strawberry 0.74 0.05 0.25 1.08 2.23 0.25
CBF 6.07 - - 6.71 6.31 5.13
CinCECGTorso 31.76 - - 38.69 30.94 33.93
TwoPatterns 6.56 - - 7.76 5.6 5.75
ECG5000 7.73 - - 7.81 10.06 7.56
Plane 3.8 - - 4.71 7.35 3.1
FacesUCR 5.36 - - 6.92 7.89 4.98
NI-ECG2 2.43 - - 3.16 7.51 1.53
Average Rank 3.27 1.0 3.0 3.93 4.4 2.07
@@
Plausibility results are more scattered, particularly for multivariate datasets, where the variability across datasets and outlier detection models leads to average ranks ranging between 2 and 3 for all methods. Under these conditions, Multi-SpaCE achieves second-best solutions in many datasets and models. It ranks second overall when using Autoencoders to assess plausibility and achieves the best results when Local Outlier Factors are used.
@@
AWR 0.78 0.74 0.67 0.7 0.53 0.49 0.46 0.49 0.35 0.32 0.27 0.3
BasicMotions 0.49 0.4 0.43 0.34 0.39 0.36 0.27 0.28 0.22 0.14 0.06 0.06
Cricket 0.75 0.74 - 0.76 0.4 0.35 - 0.33 0.25 0.18 - 0.14
Epilepsy 0.48 0.59 0.58 0.61 0.24 0.31 0.3 0.34 0.07 0.12 0.14 0.16
NATOPS 0.79 0.62 0.71 0.65 0.31 0.25 0.26 0.28 0.24 0.19 0.21 0.21
PEMS-SF 0.54 0.5 - 0.51 0.23 0.19 - 0.22 0.06 0.08 - 0.05
PenDigits 0.19 0.3 0.56 0.27 0.27 0.34 0.4 0.32 0.22 0.32 0.63 0.25
RacketSports 0.74 0.74 0.75 0.73 0.35 0.38 0.36 0.36 0.13 0.14 0.14 0.14
SR-SCP1 0.62 0.42 0.61 0.44 0.19 0.2 0.19 0.21 0.09 0.1 0.16 0.11
UWave 0.45 0.72 0.72 0.62 0.37 0.43 0.46 0.48 0.39 0.42 0.4 0.42
Average Rank 2.6 2.0 2.88 2.1 2.3 2.4 2.0 2.6 2.2 2.4 2.38 2.1
@@
Table 4: Plausibility (OS) results in Multivariate datasets.
For univariate datasets, the results are clearer. GlacierAE performs best when Autoencoders and Local Outlier Factors are used to asses plausibility, while Multi-SpaCE achieves the third-best average rank across models, positioning it in the middle of the performance spectrum across different scenarios. However, let us remark again that most of the methods beating Multi-SpaCE in some datasets (e.g., Glacier, GlacierAE, DiscoX and AB-CF) often achieve very low validity scores. These results indicate Multi-SpaCE’s ability to generate plausible counterfactuals across various scenarios while maintaining perfect validity, unlike most of the competing approaches.
Dataset OS(AE) OS(IF)
Coffee 0.98 0.98 0.32 0.98 0.96 0.98 0.32 0.42 1.32 0.31 0.4 0.37
ECG200 0.69 0.84 0.39 0.85 1.02 0.86 0.33 0.34 0.18 0.32 0.4 0.37
FordA 0.35 0.39 0.1 0.38 0.4 0.47 0.47 0.54 0.36 0.53 0.12 0.55
Gunpoint 0.77 0.78 0.56 0.81 0.86 0.79 0.2 0.19 0.32 0.2 0.25 0.2
HandOutlines 0.29 0.95 0.35 0.33 0.48 0.48 0.12 0.11 0.11 0.12 0.17 0.12
ItalyPower 0.61 0.63 0.27 0.69 0.97 0.65 0.24 0.32 0.26 0.25 0.3 0.24
PPOC 0.14 0.13 0.03 0.08 0.52 0.16 0.16 0.19 0.07 0.1 0.33 0.2
Strawberry 0.35 0.28 0.18 0.41 0.74 0.32 0.15 0.16 0.13 0.12 0.21 0.15
CBF 0.93 - - 0.99 0.87 0.93 0.42 - - 0.51 0.45 0.46
CinCECGTorso 0.86 - - 0.75 0.6 0.77 0.42 - - 0.4 0.33 0.37
TwoPatterns 0.74 - - 0.75 1.03 0.65 0.36 - - 0.37 0.5 0.35
ECG5000 0.39 - - 0.54 0.55 0.38 0.32 - - 0.43 0.36 0.26
Plane 0.5 - - 0.79 1.09 0.54 0.24 - - 0.28 0.46 0.25
FacesUCR 0.53 - - 0.6 0.65 0.55 0.43 - - 0.47 0.46 0.45
NI-ECG2 0.1 - - 0.1 0.13 0.12 0.18 - - 0.17 0.29 0.2
Average Rank 2.2 3.38 1.25 3.2 4.2 3.2 2.2 3.88 2.88 2.67 4.0 2.8
Dataset OS(LOF)
NG Glacier Glacier(AE) AB-CF DiscoX Multi-SpaCE
Coffee 0.29 0.38 26.48 0.26 4.86 0.4
ECG200 0.26 0.2 0.18 0.47 0.94 0.34
FordA 0.26 0.35 0.3 0.27 0.25 0.4
Gunpoint 0.1 0.09 0.08 0.09 0.16 0.1
HandOutlines 0.03 0.02 0.02 0.03 0.11 0.03
ItalyPower 0.11 0.19 0.11 0.1 0.14 0.11
PPOC 0.18 0.07 0.04 0.04 1.17 0.1
Strawberry 0.11 0.11 0.11 0.1 0.6 0.11
CBF 0.16 - - 0.14 0.14 0.14
CinCECGTorso 4.25 - - 1.92 1.09 1.29
TwoPatterns 0.29 - - 0.31 0.59 0.33
ECG5000 0.23 - - 0.34 0.43 0.25
Plane 0.07 - - 0.13 0.37 0.09
FacesUCR 0.27 - - 0.29 0.28 0.27
NI-ECG2 0.02 - - 0.03 0.13 0.03
Average Rank 2.4 3.0 2.25 2.33 4.13 2.8
Table 5: Plausibility (OS) results in univariate datasets.
@@
Sparsity and contiguity are closely related properties. Methods that restrict solutions to one subsequence per channel, such as NG and COMTE, excel in contiguity. However, these approaches often result in unnecessary changes, reflected in worse sparsity scores (see Appendix C). Therefore, sparsity and the number of sub-sequences should be measured jointly. To assess the different methods following this intuition, we simply compute the arithmetic mean between the normalized sparsity and the normalized number of subsequences. Ideally, the metric value should be close to 0, indicating both low sparsity and a minimal number of subsequences. Figure 2 presents the results as boxplots. For univariate datasets, Multi-SpaCE achieves the lowest median value, consistently outperforming other methods. In multivariate datasets, Multi-SpaCE achieves the best average rank across most datasets, with a few exceptions like COMTE in PEMS-SF and DiscoX in PenDigits and ArticularyWordRecognition. These results demonstrate Multi-SpaCE’s ability to balance sparsity and contiguity effectively across diverse datasets.

### v2

@@ 4.4 Evaluation metrics
  • Proximity: quantified with the $\ell_{2}$ distance between the original sample $\mathbf{x}$ and the counterfactual $\mathbf{x^{\prime}}$.
  • Sparsity: evaluated as the number of changes of the counterfactual, normalized by the total length of the time series: $\frac{||\mathbf{M}||_{0}}{L\times C}$. A lower score is desirable.
@@
We also compare the execution times and present them in the Supplementary Material.
@@ 4.5 Results
In this section, we present the results in terms of validity, proximity, sparsity, plausibility, and contiguity. For all metrics (except for validity), reported results are calculated exclusively on valid CFEs. The results are organized into tables specific to each metric, with separate subtables for univariate and multivariate datasets. The best performances for each dataset are highlighted in bold, the second-best results are underlined, and dashes indicate that no valid solutions were provided for that method/dataset pair. Additionally, we rank the methods for each dataset and metric and report their average rank to provide an intuitive performance comparison across different settings.
@@
AWR 0.97 0.4 0.09 1.0
BasicMotions 1.0 0.22 0.08 1.0
Cricket 1.0 0.44 - 1.0
Epilepsy 1.0 0.38 0.55 1.0
NATOPS 0.9 0.67 0.11 1.0
PEMS-SF 0.99 0.77 - 1.0
PenDigits 1.0 0.76 0.88 1.0
RacketSports 0.98 0.66 0.25 1.0
SR-SCP1 1.0 0.59 0.55 1.0
UWave 0.92 0.75 0.38 1.0
Average Rank 1.5 3.2 3.75 1.0
@@
Coffee 1.0 0.75 0.46 1.0 1.0 1.0
ECG200 1.0 0.17 0.13 0.84 0.86 1.0
FordA 1.0 0.2 0.38 0.98 0.71 1.0
Gunpoint 1.0 0.38 0.3 1.0 0.91 1.0
HandOutlines 1.0 1.0 0.99 0.87 0.94 1.0
ItalyPower 1.0 0.06 0.13 0.48 0.39 1.0
PPOC 1.0 0.99 0.57 0.97 0.92 1.0
Strawberry 1.0 0.88 0.62 1.0 0.83 1.0
CBF 1.0 - - 1.0 1.0 1.0
CinCECGTorso 1.0 - - 1.0 1.0 1.0
TwoPatterns 1.0 - - 1.0 1.0 1.0
ECG5000 1.0 - - 0.67 0.76 1.0
Plane 1.0 - - 1.0 1.0 1.0
FacesUCR 1.0 - - 1.0 1.0 1.0
NI-ECG2 1.0 - - 1.0 1.0 1.0
Average Rank 1.0 4.38 5.5 2.2 2.67 1.0
@@
Proximity results are presented in Table 3. While Multi-SpaCE does not directly optimize for proximity, it achieves competitive results. For multivariate datasets, Multi-SpaCE achieves the best average rank by a small margin. In univariate datasets, Glacier is the top performer overall, frequently achieving the lowest proximity scores, followed by Multi-SpaCE ranking as the second-best method on average, and performing the best on most datasets where Glacier is not applicable. However, note that the number of valid counterfactuals generated by Glacier is often very low, thus reducing the relevance of its good proximity score. For example, for ECG200, Glacier obtains a proximity score of 0.42 on valid counterfactuals, but 83% of the generated counterfactuals are not valid.
@@
AWR 25.8 39.43 10.18 16.52
BasicMotions 109.12 125.96 53.91 66.71
Cricket 71.0 98.58 - 51.67
Epilepsy 20.85 21.58 14.08 13.04
NATOPS 8.75 18.77 8.9 8.38
PEMS-SF 4.52 11.93 - 5.52
PenDigits 95.11 102.05 82.73 43.95
RacketSports 65.27 73.45 38.91 50.97
SR-SCP1 627.08 1127.41 356.33 285.74
UWave 34.18 25.97 17.42 15.2
Average Rank 2.7 3.7 1.75 1.4
@@
Coffee 1.32 0.37 12.32 1.33 2.93 1.2
ECG200 3.04 0.42 1.97 3.38 4.91 2.56
FordA 13.66 0.59 5.6 16.43 12.45 7.29
Gunpoint 2.31 0.43 4.02 3.07 4.1 2.27
HandOutlines 2.45 0.14 0.31 3.44 6.87 0.57
ItalyPower 1.65 0.39 1.28 1.52 1.95 1.53
PPOC 0.33 0.07 0.25 0.37 1.13 0.19
Strawberry 0.81 0.11 0.22 0.8 1.81 0.37
CBF 5.85 - - 7.67 5.71 5.95
CinCECGTorso 30.55 - - 30.77 26.7 18.71
TwoPatterns 6.46 - - 8.84 5.45 6.02
ECG5000 5.02 - - 7.56 9.74 6.08
Plane 3.86 - - 5.83 6.7 3.56
FacesUCR 5.7 - - 7.6 9.51 5.47
NI-ECG2 2.21 - - 3.67 10.59 1.83
Average Rank 3.2 1.0 3.0 4.0 4.33 2.2
@@
The plausibility results exhibit greater variability, particularly for univariate datasets (see Table 4), where no single method consistently outperforms others across all datasets and outlier detection models. Notably, Glacier(AE) achieves the highest plausibility scores when evaluated using the Autoencoder and Local Outlier Factor (LOF), but its performance drops to fourth place under Isolation Forest evaluation. Multi-SpaCE, maintains mid-range positions across evaluation models: it ranks second under Isolation Forest, fourth under the Autoencoder, and fifth under Local Outlier Factor. However, let us remark again that most of the methods beating Multi-SpaCE in some datasets (e.g., Glacier, GlacierAE, DiscoX and AB-CF) often achieve very low validity scores. These results indicate Multi-SpaCE’s ability to generate plausible counterfactuals across various scenarios while maintaining perfect validity, unlike most of the competing approaches.
Dataset OS(AE) OS(IF)
Coffee 0.99 0.98 0.32 0.98 0.96 0.98 0.37 0.41 1.32 0.3 0.44 0.4
ECG200 0.81 0.8 0.45 0.92 1.09 0.88 0.38 0.44 0.4 0.45 0.49 0.39
FordA 0.35 0.31 0.1 0.37 0.4 0.47 0.52 0.5 0.36 0.63 0.08 0.58
Gunpoint 0.81 0.76 0.58 0.78 0.85 0.8 0.21 0.16 0.35 0.2 0.29 0.21
HandOutlines 0.27 0.26 0.26 0.32 0.48 0.3 0.12 0.12 0.12 0.11 0.17 0.12
ItalyPower 0.55 0.77 0.26 0.62 0.96 0.61 0.23 0.41 0.22 0.23 0.32 0.22
PPOC 0.11 0.12 0.03 0.12 0.38 0.12 0.16 0.19 0.11 0.16 0.28 0.18
Strawberry 0.32 0.29 0.15 0.33 0.75 0.36 0.13 0.14 0.08 0.12 0.18 0.15
CBF 0.96 - - 0.95 0.87 0.9 0.5 - - 0.47 0.52 0.43
CinCECGTorso 0.79 - - 0.82 0.63 0.8 0.41 - - 0.47 0.31 0.4
TwoPatterns 0.7 - - 0.76 1.03 0.67 0.38 - - 0.34 0.51 0.34
ECG5000 0.48 - - 0.42 0.57 0.4 0.28 - - 0.37 0.36 0.26
Plane 0.51 - - 0.74 1.04 0.55 0.26 - - 0.26 0.45 0.26
FacesUCR 0.51 - - 0.61 0.68 0.53 0.41 - - 0.48 0.52 0.41
NI-ECG2 0.1 - - 0.1 0.15 0.11 0.19 - - 0.19 0.33 0.2
Average Rank 2.73 2.5 1.0 3.33 4.33 3.07 2.27 3.62 2.75 2.53 4.27 2.33
Dataset OS(LOF)
NG Glacier Glacier(AE) AB-CF DiscoX Multi-SpaCE
Coffee 0.44 0.38 27.42 0.3 3.99 0.46
ECG200 0.44 0.31 0.2 0.5 1.0 0.49
FordA 0.3 0.3 0.29 0.29 0.25 0.39
Gunpoint 0.11 0.08 0.11 0.1 0.2 0.11
HandOutlines 0.03 0.03 0.03 0.03 0.1 0.03
ItalyPower 0.11 0.3 0.08 0.09 0.14 0.1
PPOC 0.06 0.06 0.04 0.06 0.53 0.07
Strawberry 0.11 0.09 0.08 0.09 0.37 0.11
CBF 0.18 - - 0.13 0.2 0.13
CinCECGTorso 1.49 - - 2.68 0.81 0.74
TwoPatterns 0.34 - - 0.26 0.6 0.3
ECG5000 0.24 - - 0.26 0.48 0.25
Plane 0.07 - - 0.12 0.3 0.07
FacesUCR 0.24 - - 0.31 0.4 0.23
NI-ECG2 0.02 - - 0.03 0.2 0.02
Average Rank 2.53 2.5 2.0 2.33 4.47 2.6
Table 4: Outlier Score (OS) results in univariate datasets. The lower the OS the higher the plausibility.
In the case of multivariate datasets (see Table 5), the results are more conclusive. Multi-SpaCE demonstrates superior performance across evaluation models, achieving the lowest average outlier scores under AE and IF, and second-best results under LOF. This consistency is driven by its ability to produce counterfactuals that are ranked first or second across nearly all datasets, highlighting its robustness in generating plausible instances while simultaneously preserving high performance in other evaluation dimensions. COMTE typically ranks second in plausibility, particularly when plausibility is assessed using Autoencoder and Local Outlier Factor models. These results indicate Multi-SpaCE’s ability to generate plausible counterfactuals across various scenarios while maintaining perfect validity, unlike most of the competing approaches.
@@
AWR 0.8 0.8 0.65 0.66 0.5 0.55 0.32 0.48 0.32 0.43 0.16 0.3
BasicMotions 0.51 0.71 0.23 0.33 0.38 0.7 0.1 0.28 0.16 0.54 0.0 0.05
Cricket 0.71 0.78 - 0.71 0.31 0.43 - 0.29 0.12 0.14 - 0.13
Epilepsy 0.48 0.56 0.64 0.56 0.25 0.29 0.31 0.28 0.1 0.1 0.12 0.11
NATOPS 0.82 0.72 0.74 0.61 0.3 0.36 0.27 0.29 0.24 0.39 0.12 0.2
PEMS-SF 0.52 0.54 - 0.51 0.23 0.23 - 0.22 0.06 0.12 - 0.05
PenDigits 0.2 0.4 0.56 0.26 0.29 0.35 0.41 0.32 0.22 0.42 0.66 0.23
RacketSports 0.72 0.53 0.78 0.69 0.35 0.2 0.37 0.33 0.12 0.07 0.16 0.12
SR-SCP1 0.61 0.42 0.49 0.45 0.2 0.21 0.19 0.22 0.11 0.07 0.12 0.13
UWave 0.49 0.72 0.75 0.61 0.44 0.5 0.44 0.44 0.37 0.46 0.36 0.4
Average Rank 2.3 2.5 3.0 1.7 2.1 3.1 2.12 1.9 2.0 2.8 2.38 2.3
@@
Table 5: Outlier Score (OS) results in Multivariate datasets. The lower the OS the higher the plausibility.
@@
Sparsity and contiguity are closely related properties. Methods that restrict solutions to one subsequence per channel, such as NG and COMTE, excel in contiguity. However, these approaches often result in unnecessary changes, reflected in worse sparsity scores (see Supplementary Material). Therefore, sparsity and the number of sub-sequences should be measured jointly. To assess the different methods following this intuition, we simply compute the arithmetic mean between the normalized sparsity and the normalized number of subsequences. Ideally, the metric value should be close to 0, indicating both low sparsity and a minimal number of subsequences. Figure 2 presents the results as boxplots. For univariate datasets, Multi-SpaCE achieves the lowest median value, consistently outperforming other methods. In multivariate datasets, Multi-SpaCE achieves the best median value, with a few exceptions like COMTE in PEMS-SF and DiscoX in BasicMotions and Epilepsy. These results demonstrate Multi-SpaCE’s ability to balance sparsity and contiguity effectively across diverse datasets.
@@
Figure 3 presents illustrative examples of counterfactual explanations generated by Multi-SpaCE and baseline methods. AB-CF typically produces counterfactuals with a large number of changes but a limited number of subsequences. COMTE, by design, modifies entire channels, while DiscoX tends to generate highly sparse counterfactuals, but with low probability of achieving valid solutions. In contrast, Multi-SpaCE produces counterfactuals that modify a limited number of subsequences, often resulting in smoother and more interpretable transitions compared to the abrupt changes on the original signal observed in methods like DiscoX.
Figure 3: Examples of counterfactual explanations. The red and blue lines represent the original input and the generated counterfactual respectively. The background in red indicates the timestamps in which the counterfactual value differs from the original one. Not valid counterfactuals are likely for some methods (i.e., DiscoX has a validity of 9% in AWR).
It is important to note that the counterfactuals shown in Figure 3 correspond to the solutions with the highest utility score, computed using the fixed objective weights described in Section 4.1. However, recall that Multi-SpaCE returns a Pareto front of non-dominated solutions, allowing the end-user to select the counterfactual that best aligns with their specific interpretability or plausibility requirements. Figure 4 illustrates an example of such a Pareto front for the sparsity and plausibility objectives. As shown, the most plausible counterfactuals—those with the lowest outlier scores—tend to concentrate changes at the same time steps across all channels. As sparsity is increasingly prioritized, solutions begin to distribute changes more independently across channels, which in turn raises the outlier score and reduces overall plausibility.
Figure 4: Example of Pareto Front (sparsity vs plausibility) of counterfactual explanations produced by Multi-SpaCE for instance 135 of ArticularyWordRecognition (AWR) test set.